\begin{thebibliography}{99}

\bibitem{AGZ} J. Alvarez Montaner, R. Garcia Lopez, and S. Zarzuela
Armengou, {\it Local cohomology, arrangements of subspaces and
monomial ideals}  Adv. Math.  {\bf 174}  (2003),  no. 1, 35--56. 

\bibitem{ABZ} J. Alvarez Montaner, Alberto F.Boix, and S. Zarzuela
Armengou, {\it On some Local cohomology spectral sequence}, International Mathematics Research Notices, Volume 2020, Issue 19, October 2020, Pages 6197?293

\bibitem{BBLSZ}
%\bibitem[BBLSZ]{BBLSZ}
B. Bhatt, M. Blickle, G. Lyubeznik, A. Singh and W. Zhang, {\it Local Cohomology Modules of a Smooth $\ZZ$-Algebra have Finitely Many Associated Primes}, Inventiones Mathematicae, 197 (2014) 509-519, arXiv: 1304.4692v2 [math.AC].
%\bibitem[BS98]{BroSharp}
%Markus~P. Brodmann and Rodney~Y. Sharp.
%\newblock {\em Local cohomology: an algebraic introduction with geometric
%  applications}, volume~60 of {\em Cambridge Studies in Advanced Mathematics}.
%\newblock Cambridge University Press, Cambridge, 1998.
%\bibitem[Bur]{Bur}
%L Burch.
%\newblock Codimension and andanalytic spread, 
%\newblock {\em Proc. Cambridge Philos. Soc.} 72, 1972, 369-373.

%\bibitem{HartshorneLC}
%\bibitem[Har67]{HartshorneLC}
%Robin Hartshorne.
%\newblock {\em Local cohomology}, volume 1961 of {\em A seminar given by A.
 % Grothendieck, Harvard University, Fall}.
%\newblock Springer-Verlag, Berlin, 1967.

%\bibitem{HartshorneCD}
%\bibitem[Har68]{HartshorneCD}
%Robin Hartshorne.
%\newblock Cohomological dimension of algebraic varieties.
%\newblock {\em Ann. of Math. (2)}, 88:403--450, 1968.

\bibitem{Hu}
%\bibitem[Hu1]{Hu}
C. Huneke, {\it Problems on local cohomology, in: Free resolutions in commutative algebra and algebraic
geometry (Sundance, Utah, 1990)}, 93-108, Res. Notes Math. 2, Jones and Bartlett, Boston, MA, 1992.

%\bibitem{H-L}
%\bibitem[HL90]{H-L}
%Craig Huneke and Gennady Lyubeznik.
%\newblock On the vanishing of local cohomology modules.
%\newblock {\em Invent. Math.}, 102(1):73--93, 1990.

\bibitem{HeSc}
%\bibitem[HeSc]{HeSc}
M. Hellus and P. Schenzel, {\it On cohomologically complete intersections}, Journal of Algebra, Volume 320 (10), 3733-3748, 2008. 

\bibitem{HS}
%\bibitem[HS]{HS} 
C. Huneke and R. Sharp, {\it Bass Numbers of Local Cohomology Modules}, Transaction of American Mathematical Society, Vol. 339, No. 2, 765-779, 1993.

%\bibitem{CohDim}
%\bibitem[HNBPW18]{CohDim}
%Daniel~J. Hern\'andez, Luis N\'u{\~n}ez-Betancourt, Juan~F. P\'erez, and
%  Emily~E. Witt.
%\newblock Cohomological dimension , Lyubeznik numbers, and connectedness in mixed characteristic
%\newblock {\em J. Algebra}, 514 (2018), 442-467
%\newblock {\em Preprint}, 2016.

\bibitem{Ka}
%\bibitem[Ka]{Ka}
M. Katzman, {\it An example of an infinite set of associated primes of a local cohomology module}, J. Algebra, 252, no. 1, 161?66, 2002.

\bibitem{Kwa}
%\bibitem{Kwa}[Kwa08] 
K. Kawasaki, {\it On finiteness properties of local cohomology modules over Cohen-Macaulay local ring} Illinois Journal of Mathematics
Volume 52, Number 3, Fall 2008, Pages 727-744

\bibitem{Ly1}
%\bibitem[Ly1]{Ly1}
G. Lyubeznik, {\it Finiteness Properties of Local Cohomology modules (an application of D-modules to commutative algebra)}, Invent. Math. 113, 41-55, 1993.

\bibitem{Ly2}
%\bibitem[Ly2]{Ly2}
G. Lyubeznik, {\it Finiteness Properties of Local Cohomology Modules for Regular Local Rings of Mixed Characteristic: the Unramified Case}, Comm. Alg. 28, 5867-5882, 2000.

\bibitem{Ly3}
%\bibitem[Ly3]{Ly3}
G. Lyubeznik, {\it $F$-modules: applications to Local Cohomology and $D$-modules in Characteristic $p>0$}, J Reine Angew. Math. 491, 65-130, 1997.

\bibitem{Ly4}
%\bibitem{Ly4}[Lyu07]
G. Lyubeznik, {\it On some local cohomology modules}, Advances in Mathematics
Volume 213, Issue 2, 20 August 2007, Pages 621-643

\bibitem{Ma}
%\bibitem[Mar]{Mar}
L. Ma, \emph{Infinitely many associated primes of local cohomology modules of ramified regular local rings}, arXiv:  arXiv:2604.10964v2

\bibitem{Matsumura}
%\bibitem[Mat80]{Matsumura}
H. Matsumura
\newblock {\it Commutative Ring Theory}, Cambridge University Press, 1990. 

\bibitem{Nu3}
%\bibitem[Nu3]{Nu3}
L. N$\acute{u}$$\tilde{n}$ez-Betancourt {\it On certain rings of differentiable type and finiteness properties of local cohomology}, Journal of Algebra, Volume 379, April 2013, Pages 1-10, arXiv: 1202.3597 [math.AC].
	
%\bibitem{Ogus}
%\bibitem[Ogu73]{Ogus}
%Arthur Ogus.
%\newblock Local cohomological dimension of algebraic varieties.
%\newblock {\em Ann. of Math. (2)}, 98:327--365, 1973.

%\bibitem{P-S}
%\bibitem[PS73]{P-S}
%C.~Peskine and L.~Szpiro.
%\newblock Dimension projective finie et cohomologie locale. {A}pplications \`a
%  la d\'emonstration de conjectures de {M}. {A}uslander, {H}. {B}ass et {A}.
%  {G}rothendieck.
%\newblock {\em Inst. Hautes \'Etudes Sci. Publ. Math.}, (42):47--119, 1973.

\bibitem{Roo} 
%\bibitem{Roo}[Roo]
J.-E. Roos, {\it Sur les foncteurs d$\acute{e}$riv$\acute{e}$s de lim. Applications}, C.R. Acad. Sci. Paris, 252:3702-3704, 1961. 

\bibitem{Si}
%\bibitem[Si]{Si}
A. K. Singh, {\it p-torsion elements in local cohomology modules}, Math. Res. Lett. 7, 165-176, 2000.

\bibitem{SS}
%\bibitem[SS]{SS}
A. K. Singh and I. Swanson, {\it Associated primes of local cohomology modules and of Frobenius powers}, Int. Math. Res. Not., no. 33, 1703-1733, 2004.
\end{thebibliography}

\documentclass[10pt]{amsart}
\usepackage{amssymb,amsmath,amsthm,amsfonts,amsopn,url}
\usepackage{amscd,amssymb,amsopn,amsmath,amsthm,graphics,amsfonts,enumerate,verbatim,calc}
\usepackage[all]{xy}
\theoremstyle{plain}
\newtheorem{thm}{Theorem}[section]
\newtheorem*{mt*}{Theorem (Huneke-Lyubeznik)}
\newtheorem*{cj*}{Conjecture}
\newtheorem*{nt*}{Notations}

\newtheorem{lemma}[thm]{Lemma}

\theoremstyle{definition}

\newcommand{\ideal}[1]{\mathfrak{#1}}

\newcommand{\m}{\ideal{m}}
\newcommand{\n}{\ideal{n}}

\newcommand{\func}[1]{\mathrm{#1} \,}

\newcommand{\height}{\func{ht}}
\newcommand{\grade}{\func{grade}}
\newcommand{\ara}{\func{ara}}

\newcommand{\Ass}{\func{Ass}}

\newcommand{\ZZ}{{\mathbb Z}}

\title[]{on local cohomology modules over ramified regular local rings}

\author[]{Rajsekhar Bhattacharyya}

\address{Dinabandhu Andrews College, Garia, Kolkata 700084, India}

\email{rbhattacharyya@gmail.com}

\thanks{}

\keywords{Local Cohomology}

\subjclass[2010]{13D45}
\begin{document}

\begin{abstract}
In this paper, we show examples of local cohomology modules over ramified regular local ring, having finite set of associated primes. In doing so we consider our ramified regular local ring as Eisenstein extension of an unramified regular local ring sitting inside it. In ramified regular local ring for extended ideal (from the unramified one) set of associative primes of a local cohomology module is always finite. Using the Mayer-Vietoris spectral sequence, we construct examples of local cohomology of non extended ideal, whose set of associated primes is finite. In particular, we choose those non extended ideals whose minimal primes are extended ideals which are all cohomologically complete intersection. These examples of local cohomology modules include the cases whose associative primes also contain prime number $p$.
\end{abstract}

\maketitle
\section{introduction}

Let $R$ be a Noetherian ring and $M$ be a module over it. For an ideal $I\subset R$ and for some integer $i\geq 0$, consider local cohomology module $H^i_I (M)$ with support in the ideal $I$. In the fourth problem of \cite{Hu}, it is asked that whether local cohomology modules of Noetherian rings have finitely many associated prime ideals. There are examples given in \cite{Si}, \cite{Ka}, and \cite{SS}, which show that the set of associated primes of $i$-th local cohomology module $H^i_I (R)$ of a Noetherian ring $R$ can be infinite. However, there are several important cases where we have the finiteness of the set of associated primes of local cohomology modules and we list them below: (1) Regular rings of prime characteristic \cite{HS}, (2) regular local and affine rings of characteristic zero \cite{Ly1}, (3) unramified regular local rings of mixed characteristic \cite{Ly2} and (4) smooth algebra over $\ZZ$ \cite{BBLSZ}. These results support Lyubeznik conjecture, (\cite{Ly1}, Remark 3.7): 

\textbf{Conjecture:} Let $R$ be a regular ring and $I\subset R$ be its ideal, then for each $i\geq 0$, $i$-th local cohomology module $H^i_I (R)$ has finitely many associated prime ideals.

Recently, in (\cite{Ma}) it is shown that in ramified regular local rings, there exist examples of local cohomology modules with ifinitely many associated primes. To investigate the cause in more detail, in this paper, as a first step, we construct examples when still we can get local cohomologies with finitely many associated primes: Let $S$ be a 
complete ramified regular local ring of mixed characteristic which is an Eisenstein extension of an unramified complete regular local ring $R$ sitting inside it. Let $J$ be an ideal of $S$. Assume that minimal primes of $J$ are the extension of an ideals $I_t$ for $t=1,\ldots, n$ of $R$ such the $I_t$'s are all cohomologically complete intersection \cite{HeSc}. Then $\Ass H^i_{J} (S)$ is finite. In (\cite{Nu3}), it is shown that if the set of associated primes of a local cohomology module does not contain the prime number $p$, then the set of associated primes is finite. These examples of local cohomology modules (in this paper) include the cases whose associative primes also can contain prime number $p$.

The proof of the main theorem uses the properties of Eisenstein extension (see, for example \cite{Matsumura} for its reference) and the Mayer-Vietoris spectral sequence (see \cite{AGZ}, \cite{ABZ} and \cite{Ly4}). 

\section{the main result}

Let $(R;m)$ be a local ring of mixed characteristic $p > 0$. We say $R$ is unramified if $p\notin {\m}^2$ and it is ramified if $p\in {\m}^2$. For normal local ring $(R,\m)$, consider the extension ring defined by $S=R[X]/f(X)$ where $f(X)=X^l+ a_1X^{l-1}+\ldots +a_l$ with $a_i\in \m$ for every $i=1,\ldots,l$ and $a_l\notin {\m}^2$. This ring $(S,\n)$ is local and it is defined as an Eisenstein extension of $R$ and $f(X)$ is known as an Eisenstein polynomial (see page 228-229 of \cite{Matsumura}). 

Here we note down the following important well known properties of Eisenstein extension:

(1) An Eisenstein extension of regular local ring is regular local (see Theorem 29.8 (i) of \cite{Matsumura}) and in this context, we observe that an Eisenstein extensions of an unramified regular local ring is a ramified regular local ring.

(2) Every ramified regular local ring is an Eisenstein extension of some unramified regular local ring (see Theorem 29.8 (ii) of \cite{Matsumura}. 

(3) The Eisenstein extension mentioned in (1) and (2) are faithfully flat (apply Theorem 23.1 of \cite{Matsumura}).

In proving our theorem, we need the Mayer-Vietoris spectral sequence, which arises from te Roos complex \cite{Roo} where in \cite{AGZ}, it is stated and proved only for defining ideal of an arrangement of linear subvarieties. In \cite{Ly4} it is extened for arbitrary commutative ring and module where \cite{ABZ}, the spectral sequnce has been constructed and properties are discussed very elaborately for more general categorical objects. Although the Mayer-Vietoris spectral sequences as observed in both \cite{AGZ} and \cite{Ly4}, differ in $E_0$ and $E_1$ page, but the coincide on $E_2$ page (see 4.2 Example \cite{ABZ} for detail).  

In this paper, we need the following theorem from \cite{Ly4} regarding the Mayer-Vietoris spectral sequence. 

\begin{thm}[Theorem on MVSS]\label{MV}
Let $A$ be an arbitrary commutative ring, let $I_1,\dots,I_n\subset A$ be ideals and
let $M$ be an $A$-module. There exists a spectral sequence
$$E_1^{-p,q}=\bigoplus_{i_0<\dots<i_p}H^q_{I_{i_0}+\dots+I_{i_p}}(M)
\Longrightarrow H^{q-p}_{I_1\cap\dots\cap I_n}(M).$$
\end{thm}

In our theorem we need the condition so that the above spectaral sequence will be degenarate on $E_2$-page. In \cite{AGZ} in Theorem 1.2 it is shown the degenaracy in $E_2$-page for a defining ideal of an arrangement of linear subvarieties, when the underlying regular ring contains field. In \cite{ABZ} (see Lemma 4.15), degenaracy condition is analysed in a general setting and ideal which is cohomologically complete intersection (i.e. local cohomology module in the support of the ideal is non zero only at its height) used for the degenaracy. In \cite{Ly4} it has been mentioned that the spectral sequence will be degenarate on $E_2$ page if $A$ and $A/I_i$ are all regular (possibly arbitrary), with $M=A$ and gives reference of Theorem 1.2 of \cite{AGZ}. In our proof, we elaborate how the spectral sequence is degenate when $A$ and $A/I_i$ are regular with $M=A$. 

\begin{lemma}
For a regular local ring $(A,\o)$ of dimension $n$, if $I, J$ are ideals generated by part of regular system of parameters, or equvalently $A/I, A/J$ are regular rings, then $A/(I+J)$ is also a regular local ring. 
\end{lemma}

\begin{proof}
Consider the ring $A/{\o}^2$, where sum of images of $I$ and $J$ forms a vector subspace $W\subset A/{\o}^2$ of dimension $t$ over $A/{\o}$. Suppose images $\bar{z}_1,\ldots,\bar{z}_t$ in $A/{\o}^2$ of the elements ${z}_1,\ldots,{z}_t$ from $(I+J)$ form a basis of $W$. Extend this basis to that of $\o/\o^2$ by the images $\bar{z}_1,\ldots,\bar{z}_t, \bar{z}_{t+1},\ldots, \bar{z}_n$ of the elements ${z}_1,\ldots,{z}_t,{z}_{t+1},\ldots, {z}_n$ from $A$. Now $({z}_1,\ldots,{z}_t)A \subset I+J$ and $z_{t+1}\notin I+J$. Since ${z}_1,\ldots,{z}_n$ form regular system of parameters, both $({z}_1,\ldots,{z}_t)A$ and $({z}_1,\ldots,{z}_t, z_{t+1})A$ are primes along with $I+J$. Now $\height ({z}_1,\ldots,{z}_t)A= t$ while $\height ({z}_1,\ldots,{z}_t, z_{t+1})A= t+1$ and we have the following inclusions $({z}_1,\ldots,{z}_t)A\subset I+J \subset ({z}_1,\ldots,{z}_t, z_{t+1})A$ (where right inclusion is strict). This gives $({z}_1,\ldots,{z}_t)A=I+J$ and we conclude.
\end{proof}

Now we prove the main result of the paper.

\begin{thm}
Let $S$ be a 
complete ramified regular local ring of mixed characteristic which is an Eisenstein extension of an unramified complete regular local ring $R$ sitting inside it. Let $J$ be an ideal of $S$. Assume that minimal primes of $J$ are the extension of an ideals $I_t$ for $t=1,\ldots, n$ of $R$ such the $S/I_t S$ are all regular rings (this is always true by the lemma below). Then $\Ass H^i_{J} (S)$ is finite.
\end{thm}

\begin{proof}

The basic idea of the proof is as follows: Since $S/I_1 S,\dots, S/I_n S$ are regular rings, the Mayer-Vietoris spectral sequence will degenrate on $E_2$-page. This will imply that $H^i_{J} (S)$ has a finite filtration such that every subquotient module of the filtration is isomorphic to a local cohomology module whose set of associated primes is finite. Thus the set of associated primes of $H^i_{J} (S)$ is finite. 

We apply the above theorem on MVSS for the regular local ring $S$ and ideal $J$ with $M=S$. Set $J_{i_0\ldots i_p}=I_{i_0}S+\dots+I_{i_p}S$ and $J_{i_0\ldots\widehat{i_j}\ldots i_p}=I_{i_0} S+\dots+I_{i_{j-1}} S+I_{i_{j+1}} S+\dots+I_{i_p} S$:
$$E_1^{-p,q}=\bigoplus_{i_0<\dots<i_p}H^q_{J_{i_0\ldots i_p}}(S)
\Longrightarrow H^{q-p}_{I_1 S\cap\dots\cap I_n S}(S)= H^{q-p}_{J}(S)$$

where differentials $E_1^{-p,q}\to E_1^{-p+1,q}$ on $E_1$-page are horizontal. Consider their restriction to $H^q_{J_{i_0\ldots i_p}}(S)\subset E_1^{-p,q}$ where they take elements of it to some local cohomology module $H^{q}_{J_{i_0\ldots\widehat{i_j}\ldots,i_p}}(S)\subset E_2^{-p+1,q}$ and this map is induced by natural inclusion of $J_{i_0\ldots\widehat{i_j}\ldots i_p}\subset J_{i_0\ldots i_p}$.


Now using above lemma we observe that all the $S/J_{i_0\ldots i_p}$'s are regular for $p=0,\ldots, n-1$. So, $\grade J_{i_0\ldots i_p} = \height J_{i_0\ldots i_p} = \ara J_{i_0\ldots i_p}= s$ (say) and $H^q_{J_{i_0\ldots i_p}} (S)= 0$ if $q\neq s$. If $J_{i_0\ldots i_p}=J_{i_0\ldots\widehat{i_j}\ldots i_p}$ and $q=s$, the differential becomes the identity map, otherwise it is a zero map. Thus on the $E_2$ page, we get $E_2^{-p,q}$'s which are direct sum of local cohomology modules of the form $H^q_{J_{i_0\ldots i_p}}(S)$ where $\height J_{i_0\ldots i_p}=q$. There are differentials $E_2^{-p,q}\to E_2^{-p+2,q-1}$ on the $E_2$-page and consider their restriction to $H^q_{J_{i_0\ldots i_p}}(S)\subset E_1^{-p,q}$ where they take elements of it to some local cohomology module $H^{q-1}_{J_{i_0\ldots\widehat{i_j}\ldots,\widehat{i_k},\ldots,i_p}}(S)\subset E_2^{-p+2,q-1}$. We claim there is no nonzero map between them. 
To see this, suppose $J_{i_0\ldots i_p}= J_{i_0\ldots\widehat{i_j}\ldots,\widehat{i_k},\ldots,i_p}=J'$ (say), then atleast one of the local cohomology modules $H^q_{J'}(S)$ or $H^{q-1}_{J'}(S)$ will be a zero module. If $J_{i_0\ldots i_p}\supset J_{i_0\ldots\widehat{i_j}\ldots,\widehat{i_k},\ldots,i_p}$ where containment is proper, then by Lemma 4.5 of \cite{ABZ} (see also the Remark 4.16 of \cite{ABZ}) there will be no nonzero map. Thus we donot have any nonzero differential on $E_2$-page and the spectral sequence degenerates at $E_2$ page i.e. $E_{\infty}^{-p,q}= E_2^{-p,q}$. So $H^{q-p}_{I_1 S\cap\dots\cap I_n S}(S)= H^{q-p}_{J}(S)$ will get a filtration whose subquotient modules are isomorphic to $E_{\infty}^{-p,q}$ with fixed $q$ and where $p$ will vary. Since at each page of the spectral sequence we have only finitely many rows, we have finite filtration and it is then sufficient to show that for each $(-p,q)$-pair, the set of asssociated primes of $E_{\infty}^{-p,q}$ is finite. From \cite{Ly2} we know for every ideal $I$ of $R$, $\Ass H^i_{I}(R)$ is finite. Now Eisestein extension is flat and this implies 
$\Ass H^i_{I S}(S)$ is also finite. So for each $(-p,q)$-pair $\Ass E_{\infty}^{-p,q}$ is finite. This concludes the proof.
\end{proof}

We can improve above theorem using the following lemma.

\begin{lemma}
For any regular local ring $(A,\o)$ and for a radical ideal $I$ of it, $I$ is a cohomoogically complete intersection if and only if $A/I$ is regular.
\end{lemma}

\begin{proof}
Clearly, if $A/I$ is regular then $I$ is cohomoogically complete intersection. To see the converse observe that for $I$ if its depth, height and number of minimal generators are same then it is generated by a regular sequence of same length. So if $x_1,\ldots,x_r$ generates $I$, we need to show that they will be linearly independent modulo ${\o}^2$. If not then, there exists a relation $\sum^{r}_{i=1}a_i x_i$ is in ${\o}^2$ with atleast one $a_i$ (say $a_1$ is a unit). This implies $x_1\in (x_2,\ldots,x_r)A + {\o}^2$. Consider the local ring $B=A/(x_2,\ldots,x_r)A$ with the maximal ideal $\bar{\o}$ so that $A/I= B/(\bar{x}_1)B$. Since $I$ is radical, $B/(\bar{x}_1)B$ is reduced and hence $\bar{x}_1$ generates radical ideal in $B$. So, by construction of $B$, $(\bar{x}_1)B=\bar{\o} B$ 
But $\bar{x}_1\in\bar{\o}^2 B$ and this give $\bar{\o}^2 B= \bar{\o} B$ and by Nakayama's lemma $\bar{\o}=0$ and we get contradiction.
\end{proof}

Now we write an improved version of the main theorem which gives examples of a class of ideals in $S$ whose local coomology has finitely many associated primes.

\begin{thm}
As above, let $S$ be a ramified regular local ring which is an Eienstein extension of an unramified regular local ring $R$ sitting inside it. Then\\
(1) Extension of cohomoogically complete intersection prime ideals are cohomoogically complete intersection.\\
(2) For every ideal $J$ in $S$ if their minimal primes $I_1 S\dots I_n S$ are extension of cohomologically complete intersection ideals $I_t$ for $t=1,\ldots, n$ of $R$ then $H^{i}_{J}(S)$ has finitely many associated primes. 
\end{thm}

\begin{proof}
By above lemma and since the Eisenstein extension is faithfully flat, their extensions in $S$ is again cohomologically complete intersection. So by above theorem $H^{i}_{J}(S)$ has finitely many associated primes. 
\end{proof}

{\bf Remark}\\

As we mentioned in the introduction that we can construct local cohomology module for non extended ideal, whose set of associated primes is finite and associated primes also contain the prime $p$. To see this, we assume the notation as described above. Consider the ideals $J_1,\ldots, J_n$ of $S$ so that $S/J_j$ is regular or equivalently cohomologically complete intersection (see Lemma 2.4) for each $j=1,\ldots, n$ and atleast one of them, say $J_1$, contains prime $p$. As for example, one can take $J_1$ as maximal ideal $\n$ of $S$. Take $J= J_1\cap\dots\cap J_n$. Now Eisenstein extension is module finite and faithfully flat and in this context we assume further that the ideals $J_2,\ldots, J_n$ satisfy conditions of theorem 1 of \cite{Kwa} and for $\n= J_1$ we already have $H^{i}_{\m S}(S)= H^{i}_{\n}(S)$, for every $i\geq 0$. So, we can apply above theorem to conclude that $H^i_J (S)$ is finite and its associated primes contain $p$.

\end{document}